the sum of 3 cpnsecutives odd integers is 40 more than the smallest. What are the integers?
Answer by elima(1433)

the sum of 3 cpnsecutives odd integers is 40 more than the smallest. What are the integers?
x= 1st integer
x+2=2nd integer
x+4=3rd integer
===================
x+x+2+x+4=x+40
3x+6=x+40
3x-x=40-6
2x=34
x=17
So the smallest integer = 17
2nd = x+2=17+2=19
3rd = x+4=17+4=21
:)